A certain rifle bullet has a mass of 8.61 g. Calculate the de Broglie wavelength of the bullet traveling at 1289 miles per hour. Physical constants can be found here.
1
Expert's answer
2017-11-18T12:46:07-0500
de Broglie relation:
λ=h/mv
Where λ is wavelength (m), m is mass (kg) v is speed (m/s), h is Planck constant
(h=6.626×10-34 m2kg/s).
λ=(6.626×〖10〗^(-34) (m^2 kg)/s)/(8.61g×1kg/1000g×1289 miles/h×(1609.344 m×1h)/(1 mile×3600 s))=1.63×〖10〗^(-32) m
Answer:
1.63×10-32 m.